OpenStudy (anonymous):

In a triangle ABC, a=11, b=5 and c=30°. Use the law of cosines to determine side c.

zepdrix (zepdrix):



zepdrix (zepdrix):

\[\Large\bf\sf c^2\quad=\quad a^2+b^2-2ab \cos\mathscr C\]

zepdrix (zepdrix):

Just need to plug some stuff in, yes?

zepdrix (zepdrix):

\[\Large\bf\sf c\quad=\quad \sqrt{11^2+5^2-2(11)(5)\cos30^o}\]